OpenStudy (anonymous):

Which of the following is equivalent to the sum of 3n plus 2 for n equals 3 to 5 ?

OpenStudy (aum):

3n + 2 put n = 3: 3(3)+2 = 9+2 = 11 put n = 4: 3(4)+2 = 12 + 2 = 14 put n = 5: 3(5)+2 = 15+2 = 17 Add: 11+14+17 = ?